OpenStudy (anonymous):

Suppose you have just enough money, in coins, tot pay for a loaf of bread priced at $1.95. You have 12 coins, all quarters and dimes. Let q= the number of quarters and d=the number of dimes. What system models the given information?

OpenStudy (alexwee123):

0.25x+0.1y=1.95 x+y=12

OpenStudy (pfenn1):

q + d = 12 0.25q + 0.10d = 1.95

OpenStudy (alexwee123):

x=quarters and y=dimes

OpenStudy (anonymous):

thanks guys!

OpenStudy (pfenn1):

you're welcome

OpenStudy (alexwee123):

de nada senor